# Non-overlapping, Schwarz-type Domain Decomposition Method for Physics and Equality Constrained Artificial Neural Networks

### v1

@@ 3.4 Proposed Domain Decomposition Method
By incorporating the conditionally adaptive augmented Lagrangian method in Algorithm 1, we present our domain decomposition training procedure for learning the solution of PDEs using neural networks in a distributed computing setting, as detailed in Algorithm 2. The algorithm begins by taking inputs such as collocation points $D$, the number of subdomains $K$, the number of inner epochs $P$ for local training, and the number of outer iterations $N$ for interface communication. Each computing processor (rank), indexed by $k$, initializes its respective subdomain $\Omega_{k}$ and collocation points $D_{k}$ by partitioning the global problem. It then sets up a local model with parameters $\theta_{k}^{0}$ with random Xavier initialization scheme [37]. and initializes the interface parameters $\textbf{q}_{k}$ and interface information $\textbf{f}_{k}$. Additionally, each rank initializes Lagrange multipliers $\bm{\lambda}$ and penalty parameters $\bm{\mu}$ for managing the various constraint functions specific to its subdomain.
The training process is conducted in parallel across all ranks. Each local model is independently trained for at least $epoch_{max}$ inner epochs using the previously provided interface information. Upon completing each local training cycle or achieving convergence, the algorithm exchanges the locally optimized interface data with neighboring subdomains, including the variables’ values, fluxes and tangential derivatives based on the relationship given in Eq.  6. This exchange occurs after each cycle and is repeated for $N$ outer iterations. Our approach is distinct from the instantaneous communication and averaging of interface data at every epoch as adopted in XPINNs [20]. The independent training facilitated by inner iterations minimizes communication overhead and ensures that each subdomain model is specifically adapted to its segment of the problem, without the immediate influence of neighboring subdomains. This method enhances both the specificity and efficiency of the learning process across the distributed network.
The output from this iterative and parallel training process is a set of trained local neural network models, each finely tuned to its part of the global problem, but collectively forming a coherent solution across the entire domain. The algorithm effectively leverages the distributed architecture of modern computational systems, balancing the computational load and minimizing communication overhead, which is particularly advantageous in environments with distributed resources.
Input : Collocation points $D$, number of subdomains $K$, number of inner epochs $E$, number of outer iterations $N$
@@
2for each rank $k\in\{1,2,...,K\}$ do
3       Initialize subdomain $\Omega_{k}$ and collocation points $D_{k}$;
4       Initialize local neural network parameters $\theta_{k}^{0}$;
5       Initialize interface parameters $\textbf{q}_{k}^{0}=\textbf{1}$;
6       Initialize interface information $\textbf{f}_{k}^{0}=\textbf{0}$;
7       Initialize Lagrange multipliers $\bm{\lambda}_{k}^{0}=\textbf{1}$;
8       Initialize penalty parameters $\bm{\mu}_{k}^{0}=\textbf{1}$;
9
10
11for $n\leftarrow 1$ to $N$ do
@@ 8       Initialize penalty parameters $\bm{\mu}_{k}^{0}=\textbf{1}$;
13             for $p\leftarrow 1$ to $E$ do
14                   Update $\theta_{k}^{p}\leftarrow\text{Train}(\theta_{k}^{p-1})$ ;
15                   Update $\textbf{q}_{k}^{p}\leftarrow\text{Train}(\textbf{q}_{k}^{p-1})$ ;
16                   Check for convergence, update $\bm{\lambda}_{k}^{p},\bm{\mu}_{k}^{p}$ and apply the breaking criterion via Algorithm 1 ;
17
18
19      Update $\textbf{f}_{k}^{n}$ by exchanging interface information between neighboring ranks.
@@ 4 Numerical Experiments
We apply our domain decomposition method to solve both forward and inverse problems, illustrating its robustness in physics and equality constrained machine learning of PDEs in a distributed setting. To evaluate the accuracy of our predictions, we consider an $n$-dimensional vector of predictions $\bm{\hat{u}}\in\mathbb{R}^{n}$ and the corresponding exact values $\bm{u}$. We define the relative Euclidean (or $\mathit{l^{2}}$) norm and the infinity norm ($\mathit{l^{\infty}}$) as follows:
$\mathcal{E}_{r}(\bm{\hat{u}},\bm{u})=\frac{\|\bm{\hat{u}}-\bm{u}\|_{2}}{\|\bm{u}\|_{2}},\quad\mathcal{E}_{\infty}(\bm{\hat{u}},\bm{u})=\|\bm{\hat{u}}-\bm{u}\|_{\infty}.$ (17)
@@ 4.1 Forward Problems
Here, we focus on two types of forward problems: Poisson’s and Helmholtz equations with different scales. Among them, solving multi-scale PDE problems has been a significant challenge for PINNs due to the mean computation of the global loss terms [38, 39]. It is crucial to highlight that Poisson’s and Helmholtz equations are considered foundational in the study of domain decomposition methods [13, 14, 15]. For instance, the classical Schwarz method for Poisson’s equation necessitates overlap between subdomains for convergence, with the rate of convergence being influenced by the overlap size [15]. In the present method, we pursue a DDM with no overlap between subdomains.
4.1.1 Poisson’s equation with a low-frequency solution
@@
$\nabla^{2}u$ $=s,$ $\text{in}\quad\Omega,$ (18)
@@
where $\nabla^{2}$ is the Laplacian operator applied to the function $u$, and $s$ is a given source term. A low-frequency analytic solution is constructed that satisfies Poisson’s equation (18):
@@
Figure 2: Low-frequency Poisson’s equation: (a) predicted solution on a $5\times 5$ partitioned domain, (b) absolute point-wise error. Each subdomain is labeled as (row, column) following matrix notation. The relative $l^{2}$ norm is $5.43\times 10^{-6}$.
We employ a $5\times 5$ Cartesian decomposition and utilize a shallow feed-forward neural network consisting of three hidden layers, with each layer containing 20 neurons and tangent hyperbolic activation function in each subdomain. The local neural network models are trained for at least 50 inner epochs, after which we justify whether to terminate the current subproblem optimization. Subsequently, the interface information is exchanged between neighboring subdomains. The neural network parameters are optimized by the L-BFGS optimizer with strong Wolfe line search function, while the Adam optimizer is used for the Robin parameters.
In the global domain, we utilize $160\times 160$ randomly distributed collocation points, including residual, boundary, and interface points. The exact number of points for each subdomain is determined by rounding off calculations based on the specific domain decomposition strategy employed. In this setting, each subdomain consists of 1024 residual points and 32 points for each boundary and interface."
@@
Figure 3: Low-frequency Poisson’s Equation: Evolution of (a) PDE constraints, (b) BC constraints, (c) objective functions, (d) Lagrange multipliers for PDE constraints, (e) Lagrange multipliers for BC constraints, (f) infinity norms (g) penalty parameters for PDE constraints, (h) penalty parameters for BC constraints, (i) relative $\mathit{l^{2}}$ norms for nine subdomains randomly selected out of the total 25 subdomains resulting from a $5\times 5$ decomposition. The labels indicate the positions of these subdomains in the global domain as (row, column).
@@
For ease of presentation, we randomly select nine subdomains out of the 25 subdomains and show the evolution of physics (PDE), boundary condition (BC) constraints, and objective functions as well as the corresponding Lagrange multipliers, penalty parameters, the $\mathit{l^{\infty}}$ and relative $\mathit{l^{2}}$ norms across outer iterations in Figure 3. The trends shown in Figs. 3(a-c) indicate that the PDE, BC constraints and objective functions converge after approximately 1000 outer iterations, achieving magnitudes on the order of $10^{-6}$, $10^{-10}$ and $10^{-4}$, respectively. This behavior reflects the higher prioritization of BC constraints over PDE constraints, with the objective function, represented by the interface losses. We observe that minimization of the loss terms for subdomains lacking $\mathcal{C}_{B}$, such as subdomain $(2,2)$, exhibited relatively larger values of convergence. This phenomenon in the interior subdomains is attributed to their reliance solely on interface information from neighboring subdomains, which evolves after each communication step, particularly during the early stages of training. Despite this, the effective interface condition ensures consistency and synchronization of predictions across all subdomains. Combined with the relatively fast convergence of the physics constraint, this reinforces the overall integrity and effectiveness of the distributed solution.
@@
Figs. 3(d) and 3(e) show that the Lagrange multipliers $\bm{\lambda}$ for the PDE and BC constraints converge to approximately $10^{3}$ and $10^{5}$, respectively. This difference is attributed to the varying penalty scaling factors $\bm{\eta}$ in the penalty parameters $\bm{\mu}$, with a larger $\eta_{B}$ for the BC, allowing the models to prioritize boundary conditions over physics, as seen in Figs. 3(g) and 3(h). Although $\mu_{B}$ grow significantly, reaching values around $10^{10}$, the adaptive, gradually increasing strategy is crucial for preventing ill-conditioning, a common issue in penalty methods. This approach ensures the numerical stability and robustness of our optimization process. Eventually, as shown in Figs. 3(f) and 3(i), the infinity norms and relative $\mathit{l^{2}}$ norms stabilize around $10^{-5}$ after approximately 2000 outer iterations.
@@
Figure 4: Low-frequency Poisson’s Equation: Evolution of interface parameters (a) $\alpha$, (b) $\beta$, (c) $\gamma$, (d) the ratio $\beta/\alpha$, (e) the ratio $\gamma/\alpha$ for nine randomly selected subdomains. The labels indicate the positions of these subdomains in the global domain as (row, column).
@@
Table 1: Generalization ability of the proposed DDM demonstrated through low-frequency Poisson’s equation: The figure displays the relative $\mathit{l^{2}}$ and $\mathit{l^{\infty}}$ norms of the global domain as the number of subdomains in the domain decomposition increases. Note that the number of neurons per layer has been reduced to 10, compared to 20 in the results shown in Figure 2.
Here, we assess the generalization ability of our proposed DDM using the same low-frequency solution of Poisson’s equation example. We strategically reduce the neural network model’s complexity to 10 neurons per layer from the original 20 neurons per layer while increasing the number of subdomains in the DDM. This reduction aims to limit the network’s ability to represent intricate functions. We keep the total number of collocation points assigned to the global domain constant, while varying the number of subdomains. Table 1 presents the mean and standard deviation of the relative $\mathit{l^{2}}$ norms and $\mathit{l^{\infty}}$ norms of the global domain, over 5 different trials. From the data, it becomes evident that a $2\times 1$ domain decomposition produces the highest error, indicating a weaker representation by the neural network. In contrast, as the number of subdomains are increased, the errors diminish markedly. This improvement empirically illustrates the generalization ability of our proposed DDM and the neural network’s capacity to resolve the complexities of the solution when more focused, smaller subdomains are employed. Specifically, the transition from a $2\times 1$ to a $5\times 5$ decomposition reveals a marked improvement in both of the error metrics, underscoring the benefits of finely partitioned domain decomposition in reducing the complexity of neural network models. Our tests show that our DDM offers advantages that go beyond simply addressing and accelerating large computational problems; it also enhances prediction accuracy. However, the trend of improved accuracy with an increasing number of subdomains is not sustainable indefinitely. As the subdomains become smaller, errors due to potential overfitting resulting from the less collocation points are likely to dominate, ultimately limiting performance improvements.
@@ 4.1.2 Poisson’s equation with a multi-scale solution
PINNs perform poorly when applied to multi-scale PDE problems because global loss terms are computed on a mean basis [38, 39]. This averaging process can obscure critical details in regions characterized by complex variability, thereby hindering the effective transmission of essential information from initial and boundary conditions into the target domain. The limitation can result in an inaccurate representation of the physical phenomena [40]. To demonstrate the effectiveness of our DDM and the overall PECANN framework in multi-scale scenarios, we modify one of the frequencies in the analytical solution from Eq. 19, resulting in the following multi-scale solution:
@@
Figure 5: Multi-scale Poisson’s equation: (a) predicted solution on a $8\times 8$ partitioned domain, (b) absolute point-wise error. The relative $l^{2}$ norm is $3.665\times 10^{-3}$.
In this case, we utilize $640\times 640$ randomly distributed collocation points, in the global domain, with an $8\times 8$ decomposition. All other modeling and optimizer settings remain the same as in Figure 2. Figures 5(a,b) illustrates the distributions of the predicted solution and the absolute point-wise error for the best trial, respectively. From Fig. 5(b), the maximum error is on the order of magnitude $10^{-2}$, confined in the subdomain (1, 7). Subdomains lacking boundary data generally exhibit higher errors, which contrasts with the findings from the low-frequency case, highlighting the increased complexity of multi-scale problems.
Figure 6: Multi-scale Poisson’s Equation: Evolution of (a) PDE constraints, (b) BC constraints, (c) objective functions, (d) relative $\mathit{l^{2}}$ norms, (e) the ratio $\beta/\alpha$, (e) the ratio $\gamma/\alpha$ for nine randomly selected subdomains.
We randomly select nine out of the 64 subdomains to illustrate the evolution of the PDE and BC constraints, objective functions, relative $\mathit{l^{2}}$ norms, and two interface ratios in Fig. 6. Figs. 6(a-c) demonstrate that the PDE, BC constraints, and objective functions converge after roughly 5000 outer iterations, reaching magnitudes of $10^{-3}$, $10^{-5}$, and $10^{2}$, respectively. Compared to the previous case in Fig. 3, the larger values and extended convergence time underscore the increased difficulty of the multi-scale problem. Notably, the objective function in Fig. 6(c) spikes significantly at the start of training, indicating that the algorithm prioritizes on reducing the more challenging physics and BC constraints. The subdomain (1,7), which exhibits the maximum error in Fig. 5(b), appears to stall in a local minimum, showing slower convergence of its BC constraints. Nonetheless, the evolution of these three loss terms follows the intended prioritization strategy, resulting in a prediction with good accuracy, as shown in the Fig. 6(d).
@@
4.1.3 Helmholtz equation with a multi-scale solution

### v2

@@ 3.4 Proposed Domain Decomposition Method
Algorithm 2 presents our domain decomposition training procedure for learning the solution of PDEs in a distributed computing setting using the PECANN framework with a conditionally adaptive augmented Lagrangian method. The algorithm begins by taking inputs such as domain $\Omega$, collocation points $D$, the number of subdomains $K$, the number of inner epochs $P$ for local training, the number of outer iterations $N$ for interface communication, and the local neural network parameter $\theta$. Each computing processor (rank), indexed by $k$, initializes its respective subdomain $\Omega_{k}$ and collocation points $D_{k}$ by partitioning the global problem. It then sets up a local model with parameters $\theta_{k}^{0}$ with random Xavier initialization scheme [38]. and initializes the interface parameters $\textbf{q}_{k}$ and interface information $\textbf{f}_{k}$. Additionally, each rank initializes Lagrange multipliers $\bm{\lambda}$ and penalty parameters $\bm{\mu}$ for managing the various constraint functions specific to its subdomain.
The training process is conducted in parallel across all ranks. Each local model is independently trained for at least $epoch_{min}$ inner epochs using the previously provided interface information. At the end of each local training cycle or upon reaching convergence (i.e. line 13 in Alg. 2), the algorithm exchanges locally optimized interface data–comprising the predicted solution, fluxes, and tangential derivatives–with neighboring subdomains, following the relationship specified in Eq. 6. This exchange only takes place after each local training cycle and is repeated for $N$ outer iterations. The independent training of subdomain networks for a fixed number of inner iterations defers communication overhead to the end of the local training and ensures that each subdomain model is specifically adapted to its segment of the problem, without the immediate influence of neighboring subdomains. This method enhances both the specificity and efficiency of the learning process across the distributed network
The output from this iterative and parallel training process is a set of trained local neural network models, each finely tuned to its part of the global problem, but collectively forming a coherent solution across the entire domain.
@@
Input : Domain $\Omega$, collocation points $D$, number of subdomains $K$, number of inner epochs $P$, number of outer iterations $N$, local model parameters $\theta$
2
3for each rank $k\in\{1,2,...,K\}$ do
       $\Omega_{k}\&D_{k}\leftarrow\Omega\&D$        /* Assigning subdomain & collocation points */
       $\theta_{k}^{0}$                                /* Initializing local model parameters */
       $\textbf{q}_{k}^{0}=\textbf{1}$                             /* Initializing interface parameters */
       $\textbf{f}_{k}^{0}=\textbf{0}$                             /* Initializing interface information */
       $\bm{\lambda}_{k}^{0}=\textbf{1}$                            /* Initializing Lagrange multipliers */
       $\bm{\mu}_{k}^{0}=\textbf{1}$                            /* Initializing penalty parameters */
4
5
6for $n\leftarrow 1$ to $N$ do
@@
8             for $p\leftarrow 1$ to $P$ do
9                   $\theta_{k}^{p}\leftarrow\text{Update}(\theta_{k}^{p-1})$
10                   $\textbf{q}_{k}^{p}\leftarrow\text{Update}(\textbf{q}_{k}^{p-1})$
11
12                  if $\mathcal{L}_{k}^{p}>=\omega\mathcal{L}_{k}^{p-1}$ OR $p==P$ then
                         Update $\bm{\lambda}_{k}^{p},\bm{\mu}_{k}^{p}$             /* via lines 9-16 from Algorithm 1 */
13
14
15
      Update $\textbf{f}_{k}^{n}$                               /* via Eq. 6 */
16
@@ 4 Numerical Experiments
We apply our domain decomposition method to solve both forward and inverse PDE problems in a parallel fashion using the Message Passing Interface (MPI) standard. To evaluate the accuracy of our predictions, we consider an $n$-dimensional vector of predictions $\bm{\hat{u}}\in\mathbb{R}^{n}$ and the corresponding exact values $\bm{u}$. We define the relative Euclidean (or $\mathit{l^{2}}$) norm and the infinity norm ($\mathit{l^{\infty}}$) as follows:
$\mathcal{E}_{r}(\bm{\hat{u}},\bm{u})=\frac{\|\bm{\hat{u}}-\bm{u}\|_{2}}{\|\bm{u}\|_{2}},\quad\mathcal{E}_{\infty}(\bm{\hat{u}},\bm{u})=\|\bm{\hat{u}}-\bm{u}\|_{\infty},$ (17)
@@ 4.1 Forward Problems
Here, we focus on two types of forward problems: Poisson’s and Helmholtz equations. Among them, solving PDE problems with solutions that exhibit multi-scale and high wavenumber features has been a significant challenge for PINNs due to the mean computation of the global loss terms [39, 40]. It is crucial to highlight that Poisson’s and Helmholtz equations are considered foundational in the study of domain decomposition methods [13, 14, 15]. For instance, the classical Schwarz method for Poisson’s equation necessitates overlap between subdomains for convergence, with the rate of convergence being influenced by the overlap size [15]. In the present method, we pursue a DDM with no overlap between subdomains and aim to tackle both types of equations with a common framework.
4.1.1 Poisson’s equation with a low-wavenumber solution
@@
$-\Delta u$ $=s,$ $\text{in}\quad\Omega,$ (18)
@@
An exact solution that satisfies the Poisson’s equation (18) can be assumed as follows:
@@
Figure 2: Poisson’s equation with a low-wavenumber solution: (a) predicted solution on a $5\times 5$ partitioned domain, (b) absolute point-wise error. Each subdomain is labeled as (row, column) following matrix notation. The relative $l^{2}$ norm is $5.43\times 10^{-6}$.
We employ a $5\times 5$ Cartesian decomposition and utilize a shallow feed-forward neural network consisting of three hidden layers, with each layer containing 20 neurons and tangent hyperbolic activation function in each subdomain. The local neural network models are trained for a minimum of 50 inner epochs, capped at 100, after which we justify whether to terminate the current subproblem optimization. Subsequently, the interface information is exchanged between neighboring subdomains. The neural network parameters are optimized by the L-BFGS optimizer with strong Wolfe line search function, while the Adam optimizer is used for the Robin parameters.
In the global domain, we utilize $160\times 160$ randomly distributed collocation points, including residual, boundary, and interface points. The exact number of points for each subdomain is determined by rounding off calculations based on the specific domain decomposition strategy employed. In this setting, each subdomain consists of 1024 residual points and 32 points for each boundary and interface.
@@
Figure 3: Poisson’s equation with a low-wavenumber solution: Evolution of (a) PDE constraints, (b) boundary condition constraints, (c) objective functions, (d) Lagrange multipliers for PDE constraints, (e) Lagrange multipliers for boundary condition constraints, (f) infinity norms (g) penalty parameters for PDE constraints, (h) penalty parameters for boundary condition constraints, (i) relative $\mathit{l^{2}}$ norms for nine subdomains randomly selected out of the total 25 subdomains resulting from a $5\times 5$ decomposition. The labels indicate the positions of these subdomains in the global domain as (row, column).
@@
For ease of presentation, we randomly select nine subdomains out of the 25 subdomains and show the evolution of physics (PDE), boundary condition (BC) constraints, and objective functions as well as the corresponding Lagrange multipliers, penalty parameters, the $\mathit{l^{\infty}}$ and relative $\mathit{l^{2}}$ norms across outer iterations in Figure 3. The trends shown in Figs. 3(a-b) indicate that the constraints for PDE and BC are satisfied after approximately 1000 outer iterations, achieving magnitudes on the order of $10^{-6}$, $10^{-10}$, respectively. At the same time, the objective functions drop to a satisfactory level on the order of $10^{-4}$, as shown in Fig. 3(c) This behavior reflects the higher prioritization of boundary condition constraints over PDE constraints, with the objective function, represented by the interface losses. We observe that minimization of the loss terms for subdomains lacking $\mathcal{C}_{B}$, such as subdomain $(2,2)$, exhibited relatively larger values of convergence. This phenomenon in the interior subdomains is attributed to their reliance solely on interface information from neighboring subdomains, which evolves after each communication step, particularly during the early stages of training. Despite this, the effective interface condition ensures consistency and synchronization of predictions across all subdomains. Combined with the relatively rapid satisfaction of the physics constraint, this reinforces the overall integrity and effectiveness of the distributed solution.
@@
Figs. 3(d) and 3(e) show that the Lagrange multipliers $\bm{\lambda}$ for the PDE and boundary condition constraints converge to approximately $10^{3}$ and $10^{5}$, respectively. This difference is attributed to the varying penalty scaling factors $\bm{\eta}$ in the penalty parameters $\bm{\mu}$, with a larger $\eta_{B}$ for the BC, allowing the models to prioritize boundary conditions over physics, as seen in Figs. 3(g) and 3(h). Although $\mu_{B}$ grow significantly, reaching values around $10^{10}$, the adaptive, gradually increasing strategy is crucial for preventing ill-conditioning, a common issue in penalty methods. This approach ensures the numerical stability and robustness of our optimization process. Eventually, as shown in Figs. 3(f) and 3(i), the infinity norms and relative $\mathit{l^{2}}$ norms stabilize around $10^{-5}$ after approximately 2000 outer iterations.
@@
Figure 4: Poisson’s equation with a low-wavenumber solution: Evolution of interface parameters (a) $\alpha$, (b) $\beta$, (c) $\gamma$, (d) the ratio $\beta/\alpha$, (e) the ratio $\gamma/\alpha$ for nine randomly selected subdomains. The labels indicate the positions of these subdomains in the global domain as (row, column).
@@
Table 1: Generalization ability of the proposed DDM demonstrated through the Poisson’s equation with a low-wavenumber solution: The figure displays the relative $\mathit{l^{2}}$ and $\mathit{l^{\infty}}$ norms of the global domain as the number of subdomains in the domain decomposition increases. Note that the number of neurons per layer has been reduced to 10, compared to 20 in the results shown in Figure 2.
Here, we assess the generalization ability of our proposed DDM using the same low-wavenumber solution of Poisson’s equation example. We strategically reduce the neural network model’s complexity to 10 neurons per layer from the original 20 neurons per layer while increasing the number of subdomains in the DDM. This reduction aims to limit the network’s ability to represent intricate functions. We keep the total number of collocation points assigned to the global domain constant, while varying the number of subdomains. Table 1 presents the mean and standard deviation of the relative $\mathit{l^{2}}$ norms and $\mathit{l^{\infty}}$ norms of the global domain, over 5 different trials. From the data, it becomes evident that a $2\times 1$ domain decomposition produces the highest error, indicating a weaker representation by the neural network. In contrast, as the number of subdomains are increased, the errors diminish markedly. This improvement empirically illustrates the generalization ability of our proposed DDM and the neural network’s capacity to resolve the complexities of the solution when more focused, smaller subdomains are employed. Specifically, the transition from a $2\times 1$ to a $5\times 5$ decomposition reveals a marked improvement in both of the error metrics, underscoring the benefits of finely partitioned domain decomposition in reducing the complexity of neural network models. Our tests show that our DDM offers advantages that go beyond simply addressing and accelerating large computational problems; it also enhances prediction accuracy. However, the trend of improved accuracy with an increasing number of subdomains is not sustainable indefinitely. As the subdomains become smaller, errors due to potential overfitting resulting from the less collocation points are likely to dominate, ultimately limiting performance improvements.
@@ 4.1.2 Poisson’s equation with a multi-scale solution
PINNs perform poorly when applied to PDE problems with multi-scale solutions because global loss terms are computed on a mean basis [39, 40]. This averaging process can obscure critical details in regions characterized by complex variability, thereby hindering the effective transmission of essential information from initial and boundary conditions into the target domain. The limitation can result in an inaccurate representation of the physical phenomena [41]. To demonstrate the effectiveness of our DDM and the overall PECANN framework in problems with multi-scale solutions, we amplify one of the wavenumbers in the exact solution from Eq. 19, resulting in the following multi-scale solution:
@@
Figure 5: Poisson’s equation with a multi-scale solution: (a) predicted solution on a $8\times 8$ partitioned domain, (b) absolute point-wise error. The relative $l^{2}$ norm is $3.665\times 10^{-3}$.
In this case, we utilize $640\times 640$ randomly distributed collocation points, in the global domain, with an $8\times 8$ decomposition. All other modeling and optimizer settings remain the same as in Figure 2. Figs. 5(a,b) illustrates the distributions of the predicted solution and the absolute point-wise error for the best trial, respectively. From Fig. 5(b), the maximum error is on the order of magnitude $10^{-2}$, confined in the subdomain (1, 7). Subdomains lacking boundary data generally exhibit higher errors, which contrasts with the findings from the low-wavenumber case, highlighting the increased complexity of multi-scale problems.
Figure 6: Poisson’s equation with a multi-scale solution: Evolution of (a) PDE constraints, (b) boundary condition constraints, (c) objective functions, (d) relative $\mathit{l^{2}}$ norms, (e) the ratio $\beta/\alpha$, (e) the ratio $\gamma/\alpha$ for nine randomly selected subdomains.
We randomly select nine out of the 64 subdomains to illustrate the evolution of the PDE and boundary condition constraints, objective functions, relative $\mathit{l^{2}}$ norms, and two interface ratios in Fig. 6. Figs. 6(a-c) demonstrate that after roughly 5000 outer iterations, the PDE and boundary condition constraints reach magnitudes of approximately $10^{-3}$ and $10^{-5}$, respectively, while the objective functions decrease by more than four orders of magnitude to around $10^{2}$. Compared to the previous case in Fig. 3, the larger values and extended convergence time underscore the increased difficulty of the problem with a multi-scale solution. Notably, the objective function in Fig. 6(c) spikes significantly at the start of training, indicating that the algorithm prioritizes on reducing the more challenging physics and boundary condition constraints. The subdomain (1,7), which exhibits the maximum error in Fig. 5(b), appears to stall in a local minimum, showing slower convergence of its boundary condition constraints. Nonetheless, the evolution of these three loss terms follows the intended prioritization strategy, resulting in a prediction with good accuracy, as shown in the Fig. 6(d).
@@
4.1.3 Helmholtz equation with a complex-valued solution